100642 C Exercise 5.58 Part A A vertical spring with a force constant of 350 N/m is compressed 8.0 em and a 0.35-kg ball placed on top. The How high does the ball go? Ask me
Answer #1

The elastic potential energy stored in the spring is converted to gravitational potential energy

(1/2)kx^2 = m g h

The maximum height reached is

h = kx^2 / 2mg
= (350)(0.08)^2 / 2(0.35)(9.8)
= 0.3265m
=32cm